OpenStudy (anonymous):

How would I graph -4x-4y=+1

OpenStudy (anonymous):

Thank you Callisto for answering. :)

OpenStudy (anonymous):

Isolate your y variable so that this becomes a function of x. This will help you to move your equation to what is known as point-slope form (y = mx + b) How can you get your equation to resemble this?

OpenStudy (callisto):

First put x=0 into -4x-4y=+1 And solve for y, that is the y intercept Then put y=0 into -4x-4y=+1 and solve for x, that is the x-intercept After that, find a graph paper and marking the scale Next , plot the 2 points on the graph Draw a line joining the 2 points Finally , you can get your graph :)

OpenStudy (anonymous):

uhm let's see if I got this....

OpenStudy (anonymous):

4/1 -4/1 because nothing times 4 and -4 equals 1.

OpenStudy (anonymous):

right?

OpenStudy (callisto):

Example Put x =0 into -4x-4y=+1 -4 (0) -4y = 1 0-4y =1 -4y = 1 y = -1/4 So y-intercept is at (0 , -1/4) Do you understand this?

OpenStudy (anonymous):

So 1/4 and -1/4

OpenStudy (callisto):

Can you write the coordinate of x and y intercept respectively?

OpenStudy (anonymous):